OpenStudy (anonymous):

Which system has infinite solutions? Which system has (-2, -2) as the solution? Which system has (1, 3) as the solution? Which system has (-8, 0) as the solution?

OpenStudy (anonymous):

a. x - y = -2 3x - y = 0 b.x - 4y = -8 2x - 3y = -16 c.x + y = -4 -3x + 2y = 2 d.x + y = -1 4x + 4y = -4